Thor invests some money in an account earning 8% interest, compounded continuously. After how many years will he double his investment? [Use the formla A = Pert and round answer to the nearest tenth.]
Answer by fractalier(6550)

We will use
A = Pe^rt
but the amount we are interested in is 2P (because we want to double his investment)...so we have
2P = Pe^rt
e^(.08)t = 2
Now take the ln and get
.08t = ln 2
t = ln 2/.08 = 8.66 years = 8.7 years to the nearest tenth